OpenStudy (anonymous):

can someone please explain how to do this? Multiply the equation by 4p. Explain how adding (x – h)2 to the 3 values of k verifies the number of zeros of the polynomial.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

can you post a screen shot or something? this question is somehow incomplete what is k? what is 4p?

OpenStudy (anonymous):

@misty1212 its just that

OpenStudy (misty1212):

i see there must be more to it on some other page what is "the equation"? what are "the three values of k"?

OpenStudy (misty1212):

if you multiply by \(4p\) you get \[(x-h)^2+4pk=0\]

OpenStudy (misty1212):

solving gives \[(x-h)^2=-4pk\\ x-h=\pm\sqrt{-4pk}\\ x=h\pm\sqrt{-4pk}\\ x=h\pm2\sqrt{-pk}\]

OpenStudy (anonymous):

im not sure I understand